# Global Convergence of Policy Gradient for Entropy Regularized Linear-Quadratic Control with multiplicative noise

### v1

@@ Abstract
Reinforcement Learning (RL) has emerged as a powerful framework for sequential decision-making in dynamic environments, particularly when system parameters are unknown. This paper investigates RL-based control for entropy-regularized Linear Quadratic control (LQC) problems with multiplicative noises over an infinite time horizon. First, we adapt the Regularized Policy Gradient (RPG) algorithm to stochastic optimal control settings, proving that despite the non-convexity of the problem, RPG converges globally under conditions of gradient domination and near-smoothness. Second, based on zero-order optimization approach, we introduce a novel model free RL algorithm: Sample-Based Regularized Policy Gradient (SB-RPG). SB-RPG operates without knowledge of system parameters yet still retains strong theoretical guarantees of global convergence. Our model leverages entropy regularization to accelerate convergence and address the exploration versus exploitation trade-off inherent in RL. Numerical simulations validate the theoretical results and demonstrate the efficacy of SB-RPG in unknown-parameters environments.
@@ 1 Introduction
Reinforcement Learning (RL) is a subfield of machine learning that focuses on training agents to make sequential decisions by interacting with dynamic environments. Unlike supervised learning, which relies on labeled datasets, RL agents learn through trial and error, guided by a reward signal that quantifies the desirability of their actions [1]. The ultimate goal is to discover an optimal policy—a mapping from states to actions which maximizes the cumulative long-term rewards. In recent years, RL has revolutionized the field, achieving human-level performance in domains ranging from game playing [2] to robotics [3], and autonomous driving [4].
@@
Optimal control theory seeks to design control policies that maximize a predefined performance criterion for dynamic systems. RL and optimal control are naturally aligned in their fundamental principles as both approaches incorporate decision-making considerations. However, traditional optimal control requires complete knowledge of all environmental parameters to make decisions. Obtaining such precise parameters is infeasible in real-world scenarios, making the application of optimal control particularly challenging. RL based control in situations where the system parameters are unknown has achieved significant success in recent years. Linear-quadratic (LQ) control problem, as one of the most fundamental problems in control theory, has attracted considerable attention and has been extensively studied in the RL based control literature. For example, in the continues time setting, Wang and Zhou [5] adopt the RL method to solve mean–variance portfolio problem, Wang et,al. [6] carry out a complete theoretical analysis of RL based LQ control problem, Li et al. [7] employ a policy iteration RL approach to investigate LQ mean-field control problems over an infinite horizon.
Policy Gradient [8] is a class of RL algorithm. It directly parameterizes the optimal policy and performs gradient descent on the policy, which makes it easy to implement and widely applicable. However, even for the most basic LQ control problem, policy gradient methods encounter a non-convex optimization landscape [9]. As a result, the convergence properties of policy gradient algorithms is a significant area of research, and numerous useful results are obtained in this field. Fazel et al. [9] prove the global convergence in the time homogeneous, infinite time horizon deterministic LQ problem. Based on [9], Gravell et al. [10] [11] extend the global convergence results to LQ with multiplicative noise through rigorous proofs, and Lai et al. [12] extend it to LQ with both multiplicative and additive noises. In the finite time setting, Hambly et al. [13] consider the convergence of policy gradient methods of LQ with additive noises and show a financial application. All of the aforementioned papers are generally composed of two parts. The first part consider global convergence of policy gradient under the assumption that the system parameters are known, which corresponds to the model-based RL (i.e., estimating the model parameters and learning). While the second part builds upon the first to analyze convergence under unknown parameters setting, which corresponds to the model-free RL (i.e., end-to-end). Hu et al. [14] survey several recent theoretical advances regarding the optimization landscape, global convergence properties, and sample complexity of gradient-based methods applied to different control problems, including but not limited to linear-quadratic (LQ) systems.
@@
A important topic in RL is the exploration-exploitation trade-off: the agent must balance exploiting known information to maximize immediate reward and exploring the environment by trying random actions in order to find potentially better actions and states. In most cases, exploration is highly resource-intensive, therefore, numerous solutions are proposed to address exploration-exploitation trade-off. The trade off between exploration and exploitation has been thoroughly studied for the LQR. For example, in [15] and is improved upon in [16]. In the discrete action space setting, $\epsilon$-greedy policy [17] and Botzman (Softmax) policy [18] are two effective and popular ways to balance exploration-exploitation and numerous developments have been made based on them. In addition to these, recent research has introduced entropy-regularized RL formulations. This approach explicitly integrates exploration into the optimization objective by including entropy as a regularization term, thereby imposing a trade-off weight on the entropy of the exploration strategy . Ahmed et al. [19] demonstrate that, even when the exact gradient is available, policy optimization remains challenging because of the complex geometry of the objective function. Moreover, employing policies with higher entropy can smooth the optimization landscape, facilitating connections between local optima. Neu et, al. [20] propose a general framework for entropy-regularized average-reward RL in Markov decision processes. It is noteworthy that while entropy-regularization has been quite useful it comes with the caveat that there are many other ways to promote policy exploration nor is entropy-regularization always effective as seen in [21].
@@
In the context of entropy-regularized RL formulation for LQ problems, there has also been substantial research progress; however, most of these results pertain to actor-critic methods, such as the previously mentioned [5] [6] [7]. In contrast, studies focusing on policy gradient based approaches for entropy-regularized LQ problems remain relatively limited. Michael et, al. [22] study the global linear convergence of policy gradient methods for finite-horizon continuous-time entropy-regularized LQ control problems. Guo et, al. In the discrete time setting, Guo et, al. [23] propose and analyzes two new policy gradient based RL method for entropy-regularized LQ problems: regularized policy gradient (RPG) and iterative policy optimization (IPO) and prove their fast convergence given exact model parameters(i.e., model based). However, they only consider additive noise and conducted their analysis solely in the model-based setting, which limits the applicability of their methods. Multiplicative noise models may be produce more robust policies, they are still more complex than the typical additive model which may result in slower convergence. To the best of our knowledge, addressing the case of multiplicative noise in both model-based and model-free settings still remains an open problem.
Our contribution
@@
This paper makes two fundamental contributions: First, We extend Regularized Policy Gradient (RPG) [23] method to stochastic optimal control scenarios and show that while the stochastic LQ is non-convex, the RPG still converges to a global minimum due to the property of gradient domination and almost smoothness, thereby enhancing the system’s robustness and significantly broadening its potential applications. Secondly, and more importantly, we employ a zero order optimization technique; we propose Sample Based Stochastic Regularized Policy Gradient (SB-RPG) method and rigorously prove its global convergence properties. SB-RPG does not require knowledge of the specific system parameters values. This theoretical guarantee enables our model to operate effectively in parameter-unknown scenarios, where as RPG cannot. Numerical simulations also support the theoretical results.
Notations
We adopt standard mathematical notation throughout this paper. For any matrix $Z\in\mathbb{R}^{n\times m}$, we denote $\|Z\|$ as the spectral norm of $Z$, $\|Z\|_{F}$ as the Frobenius norm of $Z$, $\sigma_{\min}(Z)$ and $\sigma_{\max}(Z)$ as the minimum and maximum singular values of $Z$, respectively, and $Z^{\top}$ as the transpose of $Z$.
Organization
@@
For the sake of brevity, only the proofs of the main theorems are presented in the main text, while all detailed proofs of the lemmas are relegated to the Appendix. The rest of sections are organized as follows. In section 2 we formulate the optimal control problem and transform it into a optimization problem. It is natural to consider first order method to cope with optimization problem so we give the explicit form of gradient with respect to optimization variables. In section 3, we consider Regularized Policy Gradient method proposed in [23] and provide the guarantee of global convergence. In section 4, we consider the case where all the parameters are unknown, we proposed Sample Based Regularized Policy gradient (SB-RPG) to cope with this situation. In section 5, we provide numerical experiments showing the effectiveness of our algorithm.
@@ 2 Formulation
In this section, we clearly formulate the stochastic optimal control problem over an infinite time horizon with a constant discounted rate $\gamma$ and derived the optimal feedback control policy in (5). Inspired by the structure of feedback control policy, we linearly parameterize our policy. By doing so, we transformed the optimal control problem into an finite dimensional (i.e., $n$-dimensional) optimization problem. Since optimization naturally involves consideration of first-order derivatives, we provide the explicit form of the first-order derivative.
Consider the following discrete time exploratory stochastic linear quadratic control system in the infinite time horizon:
$x_{t+1}$ $=(A+w_{t}^{x}C)x_{t}+(B+w_{t}^{u}D)u_{t}$ (1)
where $A,C,x_{t},w_{t}^{x}\in\mathbb{R},~B,w_{t}^{u}\in\mathbb{R}^{1\times n},~D\in\mathbb{R}^{n\times n},$ and $w_{t}^{x},w_{t}^{u}$ are white noises, which are distributed as follows
$\mathbb{E}[w^{x}_{t}]=0,~~\mathbb{E}[(w_{t}^{x})^{2}]=1$
$\mathbb{E}[w^{u}_{t}]=\mathbf{0}_{1\times n},~~\mathbb{E}[(w_{t}^{u})^{\top}w_{t}^{u}]=I_{n\times n}$
@@
$J_{\pi}(x):=\mathbb{E}_{\pi}\left.\left[\sum_{t=0}^{\infty}\gamma^{t}\left(Qx_{t}^{2}+u_{t}^{T}Ru_{t}-\tau\mathcal{H}(\pi(\cdot|x_{t}))\right)\right|x_{0}=x\right]$ (3)
@@
Theorem 2.1 (Optimal value functions and optimal control).
@@
$\quad-\frac{\tau}{2}(k+\log((2\pi)^{k}\operatorname{det}\Sigma^{*})+\gamma q+Tr(\Sigma^{*}(R+\gamma P(B^{\top}B+D^{\top}D))$
@@
$\min_{(K,\Sigma)\in\Omega}f(K,\Sigma)=\mathbb{E}_{x_{0}\sim\mathcal{D}}[C_{K,\Sigma}(x_{0})]=P_{K}\mu+q_{K,\Sigma}$ (10)
@@
$P_{K}=Q+K^{\top}RK+\gamma P_{K}(A^{2}+C^{2}+K^{\top}(B^{\top}B+D^{\top}D)K-2ABK)$ (11)
@@ 3 Global Convergence of Regularized Policy Gradient
$\lambda_{1}E_{K}^{\top}E_{K}\leq f(K,\Sigma)-f(K^{*},\Sigma^{*})\leq\lambda_{2}\nabla_{K}f^{\top}(K,\Sigma)\nabla_{K}f(K,\Sigma)+\frac{(1-\gamma)\text{Tr}[(\nabla_{\Sigma}C_{K,\Sigma}(x_{0}))^{2}]}{\sigma_{min}(R)}$ (14)
where $\lambda_{1}=\frac{\mu}{\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}$ and $\lambda_{2}=\frac{1}{\mu\sigma_{min}(R)}$
@@
$\|\nabla_{K}f(K,\Sigma)\|\leq\overline{\|\nabla_{K}f(K,\Sigma)\|}:=\frac{f(K,\Sigma)}{Q}\sqrt{\frac{f(K,\Sigma)-f(K^{*},\Sigma^{*})}{\lambda_{1}}}$
@@
$\|\nabla_{\Sigma}f(K,\Sigma)\|\leq\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}:=(1-\gamma)^{-1}\big[\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|+\frac{\tau}{2\sigma_{min}(\Sigma)}\big]$
@@
Lemma 3.3 (”Almost” smoothness of $f(K,\Sigma)$).
Fix $0<a<1$ and define $m=\frac{log(a)-a+1}{(a-1)^{2}}$, any $\Sigma$ and $\Sigma^{\prime}$ satisfies $aI\prec\Sigma\prec I$ and $aI\prec\Sigma^{\prime}\prec I$, we have,
$f(K^{\prime},\Sigma^{\prime})-f(K,\Sigma)$ $=S_{K}[(K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)+2(K^{\prime}-K)^{\top}E_{K}]$
@@
$\quad+\frac{Tr\left(((R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}\Sigma^{-1})(\Sigma^{\prime}-\Sigma)\right)}{(1-\gamma)}+\frac{\tau m}{2(1-\gamma)}Tr((\Sigma^{-1}\Sigma^{\prime}-I)^{2})$
From the above lemmas, we have established that $f(K,\Sigma)$ possesses the properties of Gradient Domination and is ”almost” smooth. These properties make it possible to prove global convergence. Now we need to show that one step update guarantees a decrease in $f(K,\Sigma)$. To this end, we first prove that the update of $\Sigma$ is bounded.
@@
Let $(K,\Sigma)\in\Omega$ be given such that $0\prec\Sigma\preceq I$. Assume $\tau\in(0,2\sigma_{min}(R))$. Fix $a\in(0,\text{min}\{\frac{\tau}{2\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|},\sigma_{min}(\Sigma)\})$. For $\eta_{1}\leq\frac{1}{\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}$ and $\eta_{2}\leq\frac{2(1-\gamma)a^{2}}{\tau}$, and $0<\phi=min\{\eta_{1}\mu\frac{\sigma_{min}{(R)}}{S_{K^{*},\Sigma^{*}}},\frac{\eta_{2}\sigma_{min}(R)}{2(1-\gamma)}\}<1$

### v3

@@ Abstract
Reinforcement Learning (RL) has emerged as a powerful framework for sequential decision-making in dynamic environments, particularly when system parameters are unknown. This paper investigates RL-based control for entropy-regularized linear-quadratic (LQ) control problems with multiplicative noises over an infinite time horizon. First, we adapt the regularized policy gradient (RPG) algorithm to stochastic optimal control settings, proving that despite the non-convexity of the problem, RPG converges globally under conditions of gradient domination and near-smoothness. Second, based on zero-order optimization approach, we introduce a novel model free RL algorithm: Sample-based regularized policy gradient (SB-RPG). SB-RPG operates without knowledge of system parameters yet still retains strong theoretical guarantees of global convergence. Our model leverages entropy regularization to address the exploration versus exploitation trade-off inherent in RL. Numerical simulations validate the theoretical results and demonstrate the efficacy of SB-RPG in unknown-parameters environments.
@@ 1 Introduction
Reinforcement learning (RL) is a subfield of machine learning that focuses on training agents to make sequential decisions by interacting with dynamic environments. Unlike supervised learning, which relies on labeled datasets, RL agents learn through trial and error, guided by a reward signal that quantifies the desirability of their actions [21]. The primary goal is to discover an optimal policy—a mapping from states to actions which maximizes the cumulative long-term rewards. In recent years, RL has achieved remarkable success, attaining human-level performance in domains such as game playing [4], robotics [14], and autonomous driving [13].
@@
Optimal control theory seeks to design control policies that maximize a predefined performance criterion for dynamic systems. RL and optimal control are naturally aligned in their fundamental principles as both approaches incorporate dynamical decision-making considerations. However, traditional optimal control methods require complete knowledge of all environmental parameters , which is often infeasible in real-world applications. RL-based control methods have thus gained prominence for scenarios with unknown system parameters. LQ control problem, as one of the most fundamental problems in the control theory, has attracted considerable attention and has been extensively studied in the RL based control literature. For example, in continues time settings, Wang and Zhou [23] adopt the RL method to solve mean–variance portfolio problem, Wang et, al. [24] carry out a complete theoretical analysis of RL based LQ control problem, Li et, al. [17, 18] employ a policy iteration RL approach to investigate LQ and mean-field LQ control problems over an infinite horizon.
Policy gradient methods [22] form a prominent class of RL algorithms that directly parameterize the policy and optimize it via gradient descent. Despite their simplicity and broad applicability, policy gradient methods face a non-convex optimization landscape even for basic LQ control problems [5]. Understanding their convergence properties is thus a central research topic. Fazel et al. [5] established global convergence for policy gradient in deterministic, infinite-horizon LQ settings. Subsequent works extended these results to LQ problems with multiplicative noise[8, 9], additive and multiplicative noise [15], and finite-horizon settings [11]. Typically, these studies first analyze global convergence under known system parameters (model-based RL), then extend to unknown parameter settings (model-free RL). Hu et al.[12] provide a comprehensive survey of recent theoretical advances in this area.
@@
A key challenge in RL is the exploration-exploitation trade-off: agents must balance exploiting known information to maximize immediate reward with exploring new actions to discover potentially better strategies. In most cases, exploration is highly resource-intensive, therefore, numerous solutions are proposed to address exploration-exploitation trade-off. The trade off between exploration and exploitation has been thoroughly studied for the LQR. For example, in [1] and is improved upon in [2]. In the discrete action space setting, $\epsilon$-greedy policy [25] and Botzman (Softmax) policy [20] are two effective and popular ways to balance exploration-exploitation and numerous developments have been made based on them. In addition to these, recent research has introduced entropy-regularized RL formulations. This approach explicitly integrates exploration into the optimization objective by including entropy as a regularization term, thereby imposing a trade-off weight on the entropy of the exploration strategy . Ahmed et al. [3] demonstrate that, even when the exact gradient is available, policy optimization remains challenging because of the complex geometry of the objective function. Moreover, employing policies with higher entropy can smooth the optimization landscape, facilitating connections between local optima. Neu et, al. [19] propose a general framework for entropy-regularized average-reward RL in Markov decision processes. It is noteworthy that while entropy-regularization has been quite useful it comes with the caveat that there are many other ways to promote policy exploration nor is entropy-regularization always effective as seen in [16].
@@
In the context of entropy-regularized RL for LQ problems, most research has focused on actor-critic methods [23, 24, 18], with relatively few studies addressing policy gradient approaches. Notably, Michael et al. [7] study the global linear convergence of policy gradient methods for finite-horizon continuous-time entropy-regularized LQ control problems. In discrete time settings, Guo et, al. [10] proposed and analyzed two new policy gradient based RL method for entropy-regularized LQ problems: regularized policy gradient (RPG) and iterative policy optimization (IPO) and proved their fast convergence given exact model parameters(i.e., model based). However, they only considered additive noises and conducted their analysis solely in the model-based setting, which limits the applicability of their methods. Multiplicative noise models may be produce more robust policies, they are still more complex than the typical additive model which may result in slower convergence. To the best of our knowledge, addressing the case of multiplicative noise in both model-based and model-free settings still remains largely unexplored.
@@
This paper makes two primary contributions: First, we extend the RPG [10] method to stochastic LQ scenarios with multiplicative noises, demonstrating that RPG achieves global convergence despite the non-convexity of the problem, thanks to gradient domination and almost-smoothness properties. This extension enhances robustness and broadens the applicability of RPG. Second, and more importantly, employing the zero order optimization technique, we propose sample based regularized policy gradient (SB-RPG) method and rigorously prove its global convergence properties. SB-RPG does not require knowledge of the specific system parameters values, enabling effective control in parameter-unknown environments where RPG cannot be directly applied. Our theoretical results are supported by numerical simulations.
@@
The rest of sections are organized as follows. In section 2 we formulate the optimal control problem and transform it into a optimization problem. It is natural to consider first order method to cope with optimization problem so we give the explicit form of gradient with respect to optimization variables. In section 3, we consider regularized policy gradient method proposed in [10] and provide the guarantee of global convergence. In section 4, we consider the case where all the parameters are unknown, we proposed sample based regularized policy gradient (SB-RPG) to cope with this situation. In section 5, we provide numerical experiments showing the effectiveness of our algorithm. For the sake of brevity, only the proofs of the main theorems are presented in the main text, while all detailed proofs of the lemmas are relegated to the Appendix.
@@ 2 Formulation
In this section, we clearly formulate the stochastic optimal control problem over an infinite time horizon with a constant discounted rate $\gamma$ and derived the optimal feedback control policy in (5). Inspired by the structure of feedback control policy, we linearly parameterize our policy. By doing so, we transformed the optimal control problem into an finite dimensional (i.e., $n$-dimensional) optimization problem. Since optimization naturally involves consideration of first-order derivatives, we provide the explicit form of the first-order derivative. We adopt standard mathematical notation throughout this paper. For any matrix $Z\in\mathbb{R}^{n\times m}$, we denote $\|Z\|$ as the spectral norm of $Z$, $\|Z\|_{F}$ as the Frobenius norm of $Z$, $\sigma_{\min}(Z)$ and $\sigma_{\max}(Z)$ as the minimum and maximum singular values of $Z$, respectively, and $Z^{\top}$ as the transpose of $Z$.
Consider the following discrete-time exploratory stochastic LQ control system in the infinite time horizon:
$x_{t+1}=Ax_{t}+Bu_{t}+Cx_{t}w_{t}^{x}+u_{t}^{\top}Dw_{t}^{u},$ (1)
where $A,C,x_{t},w_{t}^{x}\in\mathbb{R},~B\in\mathbb{R}^{1\times n},u,w_{t}^{u}\in\mathbb{R}^{n\times 1},~D\in\mathbb{R}^{n\times n},$ and $w_{t}^{x},w_{t}^{u}$ are white noises, which are distributed as follows
$\mathbb{E}[w^{x}_{t}]=0,~~\mathbb{E}[(w_{t}^{x})^{2}]=1,~~\mathbb{E}[w^{u}_{t}x_{t}^{x}]=\mathbf{0}_{n\times 1},$
$\mathbb{E}[w^{u}_{t}]=\mathbf{0}_{n\times 1},~~\mathbb{E}[(w_{t}^{u})^{\top}w_{t}^{u}]=I_{n\times n}.$
@@
$J_{\pi}(x):=\mathbb{E}_{\pi}\left.\left[\sum_{t=0}^{\infty}\gamma^{t}\left(Qx_{t}^{2}+u_{t}^{\top}Ru_{t}-\tau\mathcal{H}(\pi(\cdot|x_{t}))\right)\right|x_{0}=x\right].$ (3)
@@
Theorem 2.1 (Optimal value function and optimal control).
@@
$\quad-\frac{\tau}{2}(k+\log((2\pi)^{k}\operatorname{det}\Sigma^{*})+\gamma q+Tr(\Sigma^{*}(R+\gamma P(B^{\top}B+D^{\top}D)).$
@@
$\min_{(K,\Sigma)\in\Omega}f(K,\Sigma)=\mathbb{E}_{x_{0}\sim\mathcal{D}}[C_{K,\Sigma}(x_{0})]=P_{K}\mu+q_{K,\Sigma},$ (10)
@@
$P_{K}=Q+K^{\top}RK+\gamma P_{K}(A^{2}+C^{2}+K^{\top}(B^{\top}B+D^{\top}D)K-2ABK),$ (11)
$q_{K,\Sigma}=\frac{\text{Tr}(\Sigma(R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}(n+\log((2\pi)^{n}|\Sigma|)))}{1-\gamma}.$
@@ 3 Global Convergence of Regularized Policy Gradient
$\lambda_{1}E_{K}^{\top}E_{K}\leq f(K,\Sigma)-f(K^{*},\Sigma^{*})\leq\lambda_{2}\nabla_{K}f^{\top}(K,\Sigma)\nabla_{K}f(K,\Sigma)+\frac{(1-\gamma)\text{Tr}[(\nabla_{\Sigma}C_{K,\Sigma}(x_{0}))^{2}]}{\sigma_{min}(R)},$
where $\lambda_{1}=\frac{\mu}{\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}$ and $\lambda_{2}=\frac{1}{\mu\sigma_{min}(R)}$.
@@
$\|\nabla_{K}f(K,\Sigma)\|\leq\overline{\|\nabla_{K}f(K,\Sigma)\|}:=\frac{f(K,\Sigma)}{Q}\sqrt{\frac{f(K,\Sigma)-f(K^{*},\Sigma^{*})}{\lambda_{1}}},$
@@
$\|\nabla_{\Sigma}f(K,\Sigma)\|\leq\overline{\|\nabla_{\Sigma}f(K,\Sigma)\|}:=(1-\gamma)^{-1}\big[\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|+\frac{\tau}{2\sigma_{min}(\Sigma)}\big].$
@@
Lemma 3.3 (“Almost” smoothness of $f(K,\Sigma)$).
Fix $0<a<1$ and define $m=\frac{log(a)-a+1}{(a-1)^{2}}$, any $\Sigma$ and $\Sigma^{\prime}$ satisfies $aI\prec\Sigma\prec I$ and $aI\prec\Sigma^{\prime}\prec I$, we have
$f(K^{\prime},\Sigma^{\prime})-f(K,\Sigma)$
$=S_{K}[(K^{\prime}-K)^{\top}(R+\gamma P_{K}(B^{\top}B+D^{\top}D))(K^{\prime}-K)+2(K^{\prime}-K)^{\top}E_{K}]$
@@
$\quad+\frac{Tr\left(((R+\gamma P_{K}(B^{\top}B+D^{\top}D))-\frac{\tau}{2}\Sigma^{-1})(\Sigma^{\prime}-\Sigma)\right)}{(1-\gamma)}+\frac{\tau m}{2(1-\gamma)}Tr((\Sigma^{-1}\Sigma^{\prime}-I)^{2}).$
From the above lemmas, we have established that $f(K,\Sigma)$ possesses the properties of gradient domination and is “almost” smooth. These properties make it possible to prove global convergence. Now we need to show that one step update guarantees a decrease in $f(K,\Sigma)$. To this end, we first prove that the update of $\Sigma$ is bounded.
@@
Let $(K,\Sigma)\in\Omega$ be given such that $0\prec\Sigma\preceq I$. Assume $\tau\in(0,2\sigma_{min}(R))$. Fix $a\in(0,\text{min}\{\frac{\tau}{2\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|},\sigma_{min}(\Sigma)\})$. For $\eta_{1}\leq\frac{1}{\|R+\gamma P_{K}(B^{\top}B+D^{\top}D)\|}$ and $\eta_{2}\leq\frac{2(1-\gamma)a^{2}}{\tau}$, and $0<\phi=min\{\eta_{1}\mu\frac{\sigma_{min}{(R)}}{S_{K^{*},\Sigma^{*}}},\frac{\eta_{2}\sigma_{min}(R)}{2(1-\gamma)}\}<1,$
$f(K^{\prime},\Sigma^{\prime})-f(K^{*},\Sigma^{*})\leq(1-\phi)(f(K,\Sigma)-f(K^{*},\Sigma^{*})).$